A wave has the following properties amplitude = 0.
Answer #1

A =0.171m , T =0.508 s, v = 7.06 m/s

w = 2pi/T = 2*3.14/0.508 = 12.36 rad/s

v = w/k

7.06 = 12.36/k

k = 1.75 rad/s

y = 0.171*sin(12.36*t - 1.75*x)